Raisin Rich breakfast cereal is 10% protein. A dietician wants to raise the percentage of protein per serving to 20%. The high-protein grain that will be mixed with the cereal is 40% protein. How many grams of each should be mixed together to make a 12 grams serving of cereal with the proper percentage of protein?
Answer by jorel555(1290)

Your total is 12 grams of cereal. Let x be the amount at 40%, (12-x) be the amount at 10%. Then:
.4(x)+.1(12-x)=12(.2)
.4x+1.2-.1x=2.4
.3x=1.2
x=4
12-x=8
You need 4 gms of 40% protein cereal with 8 gms of 10% protein cereal to make 12 gms of 20% protein cereal!!!!!!!!